A metallic split ring has a target internal diameter of 50 mm, but records show that the diameters are normally distributed with mean 50 mm and standard deviation 0.07 mm. An acceptable diameter is one within the range 49.93 mm to 50.07 mm. What proportion of the output is unacceptable?

  • Round your answer to four decimal places.
Answer #1

Ans:

mean=50 mm

standard deviation=0.07

z(49.93)=(49.93-50)/0.07=-1

z(50.07)=(50.07-50)/0.07=1

P(unacceptable)=P(z<-1 or z>1)

=P(z<-1)+P(z>1)

=0.1587+0.1587

=0.3174